# LRQ: Optimizing Post-Training Quantization for Large Language Models by Learning Low-Rank Weight-Scaling Matrices

### v1

@@ Appendix E Sensitivity of Accumulated RMSE to the Number of Calibration Samples
Appendix F Ratio of the number of learnable parameters in LRQ to the number of pre-trained weights
| Llama 7B | |||
| Llama 13B | |||
| Llama 33B | |||
| Llama 65B |
Appendix G Comparison of Experimental Results before and after Per-token Asymmetric KV Cache Quantization
Table 11, 12, 13, 14, 15, and 16 show the comparison of experimental results before and after per-token asymmetric KV cache quantization. It can be easily seen that the performance difference before and after per-token asymmetric KV cache quantization is nearly inconsiderable no matter which quantization technique is chosen, as mentioned in Section 3.2. Furthermore, even without per-token asymmetric KV cache quantization, LRQ still outperforms prior state-of-the-art LLM post-training weight-activation quantization methods in most cases.
@@ Appendix G Comparison of Experimental Results before and after Per-token Asymmetric KV Cache Quantization
In Table 11 and 13, LRQ exhibits a slightly superior zero-shot performance on common sense reasoning tasks compared to FlexRound, which we believe is noteworthy since FlexRound already achieves the zero-shot performance on common sense reasoning tasks comparable to FP16 baselines. The close proximity in zero-shot performance between FlexRound and FP16 baselines on common sense reasoning tasks limits the potential for a substantial performance disparity between FlexRound and LRQ. Despite LRQ approaching the zero-shot performance of FP16 baselines more closely than FlexRound, the difference in zero-shot performance between FlexRound and LRQ cannot be anticipated to be large after all.
@@
Regarding the -bit weight quantization presented in Table 15 and 16, the adoption of a per-token asymmetric activation quantization scheme results in even naive rounding-to-nearest (RTN) performing closely to the levels of FP16 baselines on both common sense reasoning tasks and MMLU. As a result, while LRQ exhibits slightly higher accuracy compared to SmoothQuant and FlexRound for most Llama models, it can be concluded that SmoothQuant, FlexRound, and LRQ are nearly evenly matched.
In the context of -bit weight quantization as presented in Table 15, FlexRound achieves zero-shot accuracy levels comparable to FP16 baselines on common sense reasoning tasks, resulting in a relatively small zero-shot performance gap between FlexRound and LRQ, like the scenario depicted in Table 11 and 13. However, in the case of -bit weight quantization in Table 16, LRQ surpasses FlexRound by a margin ranging from to percent. Although these increments in five-shot accuracy on MMLU in Table 16 may seem modest compared to those in Table 12 and 14, we believe that the rise in five-shot accuracy by to percent on MMLU is significant. This is particularly noteworthy as it brings the five-shot accuracy gap between LRQ and FP16 baselines to less than percent on MMLU, while the corresponding gap between FlexRound and FP16 baselines remains more or less at two percent for Llama B and B.
Appendix H Implementation Details
@@
For the quantization scheme depicted in Figure 7, both FlexRound and LRQ are implemented in the experimental setting of QDrop (Wei et al., 2022) with the exception of the number of iterations for block-wise reconstruction, the batch size, and the learning rate. For all the Llama and Llama models, the number of iterations for block-wise reconstruction is set to for both FlexRound and LRQ. The learning rate and the batch size for FlexRound and LRQ are described in 17. Notice that when applying LRQ to Llama B, the key and value projection weights are quantized via not LRQ but FlexRound due to the presence of GQA (Ainslie et al., 2023) in Llama B. To obtain the experimental results in Table 1 and 3, per-token asymmetric KV cache quantization is applied after completing block-wise reconstruction for all the Transformer blocks. For both activation quantization and KV cache quantization, we employ rounding-to-nearest.
@@
Appendix I Comparison of Computation Cost to Complete the Quantization Process between SmoothQuant, FlexRound, and LRQ
| Method | Execution time (A) | Peak GPU memory usage |
| SmoothQuant | minutes | GB |
| FlexRound | hours minutes | GB |
| LRQ (Ours) | hours minutes | GB |
Appendix J Computation-Memory Trade-off at Test Time
Since only a quantization step size () and an integer matrix () are required during inference, once we obtain an integer matrix by setting to after , , , and are learned, there is no need to recompute the multiplication involving and at all. In other words, for inference, like other quantization methods, there are only a quantization step size and an integer matrix without the presence of , , , and . Therefore, the compute-memory tradeoff does not occur during inference.
@@
As SmoothQuant is orthogonal to block-wise reconstruction, one might wonder how the performance of FlexRound and LRQ would change when FlexRound and LRQ start their own learning process from the SmoothQuant baseline in place of the RTN baseline. Table 22 and 23 reveal the performance of ‘SmoothQuant (SQ) + FlexRound’ and ‘SmoothQuant (SQ) + LRQ’ on common sense reasoning benchmarks and the MMLU benchmark, respectively. Unfortunately, in most cases, SmoothQuant does not display its efficacy when combined with FlexRound and LRQ. Although SmoothQuant enhances five-shot performance of FlexRound on MMLU by almost two percent, ‘SQ + FlexRound’ still underperforms LRQ as well as ‘SQ + LRQ’ on MMLU, which implies that employing low-rank weight-scaling matrices would be a better choice than using full weight-scaling matrices with additional pre-processing like an uniform per-channel scaling transformation in SmoothQuant.

### v2

@@ Appendix D Sensitivity of Accumulated RMSE to the Number of Calibration Samples
Appendix E Detailed Experimental Results for Llama 3 8B and Mistral 7B
| Method | # Bits (W/A/KV) | PIQA | HellaSwag | WinoGrande | ARC-e | ARC-c | Average | WikiText2 |
| Llama B | ||||||||
| GPTQ | ||||||||
| AWQ | ||||||||
| QuIP | ||||||||
| FlexRound | ||||||||
| LRQ (Ours) |
| Mistral B | ||||||||||
| FlexRound | ||||||||||
| LRQ (Ours) | ||||||||||
[…]
| LRQ (Ours) |
To further justify the effectiveness of LRQ for other model families than Llama, we conduct additional experiments for Mistral 7B v0.1 (Jiang et al., 2023) on common sense reasoning tasks and WikiText2 in a low-bit per-channel weight-only quantization scheme. In Table 12, LRQ outperforms FlexRound in both -bit and -bit per-channel asymmetric weight-only quantization schemes. In particular, in a -bit per-channel asymmetric weight-only quantization scheme, LRQ surpasses FlexRound by about PPL on WikiText2 and by almost percent on common sense reasoning tasks.
Appendix F Comparison of Computation Cost to Complete the Quantization Process
| Method | Execution time (A) | Peak GPU memory usage |
| SmoothQuant | minutes | GB |
| FlexRound | hours minutes | GB |
| LRQ (Ours) | hours minutes | GB |
Additionally, we also measure the execution time and peak GPU memory usage while quantizing Llama B with -bit per-channel asymmetric weight-only quantization using calibration samples and a batch size of . Similar to Table 13, Table 14 shows that LRQ consumes more processing time but less peak GPU memory usage than FlexRound.
| Method | Execution time (A) | Peak GPU memory usage |
| FlexRound | hours minutes | GB |
| LRQ (Ours) | hours minutes | GB |
Appendix G Compression and Acceleration Effects after LRQ at Inference Time
Since only a quantization step size () and an integer matrix () are required during inference, once we obtain an integer matrix by setting to after , , , and are learned, there is no need to recompute the multiplication involving and at test time. In other words, for inference, like other uniform quantization methods including GPTQ, SmoothQuant, and AWQ, LRQ also requires only a quantization step size and an integer matrix without the presence of , , , and . Therefore, packing/unpacking techniques and acceleration kernels (Frantar et al., 2023; Lin et al., 2023; Park et al., 2024) can be applied to LRQ without additional effort.
Figure 5 and Table 15 exhibit compression and acceleration effects after LRQ with -bit and -bit per-channel weight-only uniform quantization. Regarding the compression effect, -bit uniform quantization shows x compression ratio, and -bit uniform quantization shows x compression ratio on the model size for Llama B. To verify the acceleration effect, we measure the matrix multiplication latency of Llama FFN layers across B to B on a single token generation. We utilize cuBLAS for FP baselines, while employing the LUT-GEMM kernel Park et al. (2024) for per-channel weight-only uniformly quantized models. As shown in Table 15, for Llama 2 B, the -bit per-channel weight-only quantized model shows x faster latency, and the -bit per-channel weight-only quantized model shows x faster latency than the FP cuBLAS baseline.
| Model | Method | CSR Avg. | Model Size [GB] | Latency [ms] |
| FP16 | ||||
| Llama 2 7B | LRQ (-bit) | |||
| LRQ (-bit) | ||||
| FP16 | ||||
| Llama 2 13B | LRQ (-bit) | |||
| LRQ (-bit) | ||||
| FP16 | ||||
| Llama 2 70B | LRQ (-bit) | |||
| LRQ (-bit) |
Appendix H Comparison of Experimental Results before and after Per-token Asymmetric KV Cache Quantization
Table 16, 17, 18, 20, 22, and 24 show the comparison of experimental results before and after per-token asymmetric KV cache quantization. It can be easily seen that the performance difference before and after per-token asymmetric KV cache quantization is nearly inconsiderable no matter which quantization technique is chosen, as mentioned in Section 3.2. Furthermore, even without per-token asymmetric KV cache quantization, LRQ still outperforms prior state-of-the-art LLM post-training weight-activation quantization methods in most cases.
@@ Appendix H Comparison of Experimental Results before and after Per-token Asymmetric KV Cache Quantization
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) | |||||||||
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) | |||||||||
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) |
@@
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) |
| Llama B | ||||||
| RTN | ||||||
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) | ||||||
| Llama B | ||||||
| RTN | ||||||
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) | ||||||
| Llama B | ||||||
| RTN | ||||||
@@
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) | |||||||||
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) | |||||||||
| Llama B | |||||||||
| RTN | |||||||||
| SmoothQuant | |||||||||
| FlexRound | |||||||||
| LRQ (Ours) |
@@
| Llama B | ||||||
| RTN | ||||||
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) | ||||||
| Llama B | ||||||
| RTN | ||||||
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) | ||||||
| Llama B | ||||||
| RTN | ||||||
| SmoothQuant | ||||||
| FlexRound | ||||||
| LRQ (Ours) |
In Table 16 and 18, LRQ exhibits a slightly superior zero-shot performance on common sense reasoning tasks compared to FlexRound, which we believe is noteworthy since FlexRound already achieves the zero-shot performance on common sense reasoning tasks comparable to FP16 baselines. The close proximity in zero-shot performance between FlexRound and FP16 baselines on common sense reasoning tasks limits the potential for a substantial performance disparity between FlexRound and LRQ. Despite LRQ approaching the zero-shot performance of FP16 baselines more closely than FlexRound, the difference in zero-shot performance between FlexRound and LRQ cannot be anticipated to be large after all.
@@
Regarding the -bit weight quantization presented in Table 22 and 24, the adoption of a per-token asymmetric activation quantization scheme results in even naive rounding-to-nearest (RTN) performing closely to the levels of FP16 baselines on both common sense reasoning tasks and MMLU. As a result, while LRQ exhibits slightly higher accuracy compared to SmoothQuant and FlexRound for most Llama models, it can be concluded that SmoothQuant, FlexRound, and LRQ are nearly evenly matched.
In the context of -bit weight quantization as presented in Table 22, FlexRound achieves zero-shot accuracy levels comparable to FP16 baselines on common sense reasoning tasks, resulting in a relatively small zero-shot performance gap between FlexRound and LRQ, like the scenario depicted in Table 16 and 18. However, in the case of -bit weight quantization in Table 24, LRQ surpasses FlexRound by a margin ranging from to percent. Although these increments in five-shot accuracy on MMLU in Table 24 may seem modest compared to those in Table 17 and 20, we believe that the rise in five-shot accuracy by to percent on MMLU is significant. This is particularly noteworthy as it brings the five-shot accuracy gap between LRQ and FP16 baselines to less than percent on MMLU, while the corresponding gap between FlexRound and FP16 baselines remains more or less at two percent for Llama B and B.
Appendix I Implementation Details
@@
For the quantization scheme depicted in Figure 8, both FlexRound and LRQ are implemented in the experimental setting of QDrop (Wei et al., 2022) with the exception of the number of iterations for block-wise reconstruction, the batch size, and the learning rate. For all the Llama and Llama models, the number of iterations for block-wise reconstruction is set to for both FlexRound and LRQ. The learning rate and the batch size for FlexRound and LRQ are described in 26. Notice that when applying LRQ to Llama B, the key and value projection weights are quantized via not LRQ but FlexRound due to the presence of GQA (Ainslie et al., 2023) in Llama B. To obtain the experimental results in Table 1 and 3, per-token asymmetric KV cache quantization is applied after completing block-wise reconstruction for all the Transformer blocks. For both activation quantization and KV cache quantization, we employ rounding-to-nearest.
@@
Appendix J Ratio of the number of learnable parameters in LRQ to the number of pre-trained weights
| Llama 7B | |||
| Llama 13B | |||
| Llama 33B | |||
| Llama 65B |
@@
As SmoothQuant is orthogonal to block-wise reconstruction, one might wonder how the performance of FlexRound and LRQ would change when FlexRound and LRQ start their own learning process from the SmoothQuant baseline in place of the RTN baseline. Table 31 and 32 reveal the performance of ‘SmoothQuant (SQ) + FlexRound’ and ‘SmoothQuant (SQ) + LRQ’ on common sense reasoning benchmarks and the MMLU benchmark, respectively. Unfortunately, in most cases, SmoothQuant does not display its efficacy when combined with FlexRound and LRQ. Although SmoothQuant enhances five-shot performance of FlexRound on MMLU by almost two percent, ‘SQ + FlexRound’ still underperforms LRQ as well as ‘SQ + LRQ’ on MMLU, which implies that employing low-rank weight-scaling matrices would be a better choice than using full weight-scaling matrices with additional pre-processing like an uniform per-channel scaling transformation in SmoothQuant.
Appendix M Process of in Eq. 2
Similar to the broadcasting process in Python Numpy, we add , , and .
To be more specific, let be
be
and be
Then, by the broadcasting process, can be expressed as